OpenStudy (anonymous):

how many trailing zeroes does 100! have

OpenStudy (anonymous):

2!!!

OpenStudy (anonymous):

:D

hartnn (hartnn):

for n! number of trailing zeros are = n/5 + n/5^2 +n/5^3 + ...... floor value of each term

hartnn (hartnn):

here n=100

hartnn (hartnn):

lol lolly! :P

OpenStudy (anonymous):

what? lol i did it rong? haha i thought trailing zeros were like how many zeros it had behind it...haha oop..:)

OpenStudy (anonymous):

k n wats floor value here

hartnn (hartnn):

floor value means largest integer small than the number.

OpenStudy (anonymous):

@lolly its 100! NOT 100...

OpenStudy (anonymous):

what? hah i thought u were excited bout the question..lol

hartnn (hartnn):

like floor of 3.25 is 3 floor og 101.101 is 101

hartnn (hartnn):

for 100! trailing zeros = 100/5 + 100/5^2 +100/5^3 +... = 20 + 4 + 0 +0 +.... =24 got this ?

OpenStudy (anonymous):

Yeah. thanks)

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

one more thing here v dnt tak 100/125 cz its comes in deci form .8